# Learning Temporal 3D Semantic Scene Completion via Optical Flow Guidance

### v1

@@ 5 Conclusion
Appendix Overview
This technical appendix consists of the following sections:
  • In Section A, we provide more ablation experiments to verify the effectiveness of FlowScene.
  • In Section B, we provide quantitative results from more experiments.
  • In Section C, we present more visual qualitative results of the SemanticKITTI val set.
@@
  • In Section D, we analyze the shortcomings of our method and directions for future work.
Appendix A More Ablation Studies
A.1 Ablation Study for Optical Flow Networks
Table 7 presents the performance of different optical flow networks. We compare several state-of-the-art methods, including PWC-Net [33], RAFT [36], and FlowFormer [11], along with our setting, GMFlow [43], which is highlighted in the last row. Our setting achieves the highest IoU of 45.01% and mIoU of 18.13%, outperforming all other methods in both metrics. These results suggest that GMFlow effectively captures motion cues and integrates them into the semantic scene completion task, providing superior performance over the other optical flow networks tested, with significantly fewer parameters.
A.2 Ablation Study for Backbone Networks
Table 8 examines the impact of different backbone networks on the performance of FlowScene. The study compares EfficientNetB7 [35], ResNet50 [9], and RepVit-M2.3 [37](our setting). Our method, using RepVit-M2.3, achieves the highest IoU of 45.01% and mIoU of 18.13%, surpassing both EfficientNetB7 (44.31% IoU, 17.63% mIoU) and ResNet50 (44.12% IoU, 16.98% mIoU). RepVit-M2.3, though achieving the best performance, maintains a relatively low parameter count of 22.4M. In comparison, EfficientNetB7 has a much higher parameter count of 63.8M, while ResNet50 is more parameter-efficient at 25.6M. RepVit-M2.3 offers a good balance between performance and parameter count, making it an ideal choice for our backbone network.
Method IoU(%) mIoU(%) Params(M)
PWC-Net+ [33] 43.31 17.13 8.8
RAFT [36] 44.12 17.56 5.3
FlowFormer [11] 44.33 17.74 18.2
GMFlow [43] 45.01 18.13 4.7
Table 7: Ablation study for optical flow networks.
Method IoU(%) mIoU(%) Params(M)
EfficientNetB7 [35] 44.31 17.63 63.8
ResNet50 [9] 44.12 16.98 25.6
RepVit-M2.3 [37] 45.01 18.13 22.4
Table 8: Ablation study for backbone networks.
Appendix B More Quantitative Results
@@
To provide a more thorough comparison, we provide additional quantitative results of semantic scene completion on the SemanticKITTI validation set in Table 9. The results further demonstrate the effectiveness of our approach in enhancing 3D scene perception performance. Compared with the previous state-of-the-art methods, FlowScene is superior to other HTCL [15] in semantic scene understanding, with a 1.00% increase in mIoU. In addition, compared with Symphonize [13], huge improvements are made in both occupancy and semantics. IoU and mIoU enhancement are of great significance for practical applications. It proves that we are not simply reducing a certain metric to achieve semantic scene completion.
@@
MonoScene [3] CVPR’2022 S 36.86 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25 11.08
TPVFormer [10] CVPR’2023 S 35.61 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 11.36
OccFormer[50] ICCV’2023 S 36.50 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86 13.46
Symphonize [13] CVPR’2024 S 41.92 56.37 27.58 15.28 0.95 21.64 28.68 20.44 2.54 2.82 13.89 25.72 6.60 30.87 3.52 2.24 0.00 8.40 9.57 5.76 14.89
VoxFormer-T[20] CVPR’2023 T 44.15 53.57 26.52 19.69 0.42 19.54 26.54 7.26 1.28 0.56 7.81 26.10 6.10 33.06 1.93 1.97 0.00 7.31 9.15 4.94 13.35
H2GFormer [40] AAAI’2024 T 44.69 57.00 29.37 21.74 0.34 20.51 28.21 6.80 0.95 0.91 9.32 27.44 7.80 36.26 1.15 0.10 0.00 7.98 9.88 5.81 14.29
HASSC [39] CVPR’2024 T 44.58 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10 14.74
HTCL [15] ECCV’2024 T 45.51 63.70 32.48 23.27 0.14 24.13 34.30 20.72 3.99 2.80 11.99 26.96 8.79 37.73 2.56 2.70 0.00 11.22 11.49 6.95 17.13
@@
Table 9: Quantitative results on the SemanticKITTI validation set. The best results are in Bold.
Appendix C More Visualizations Results
We show visualization examples on the Semantickitti validation set, as shown in Figure 7. From left to right are the input image, the corresponding optical flow and occlusion mask, the front view SSC, and the top view SSC. Due to the motion information brought by the optical flow, the location information of the scene objects is more accurate and the layout is more reasonable. We report the performance of more visual comparison results on the SemanticKITTI validation set in Figure 8. We compare with VoxFormer [20] and BRGScene [14]. In general, our method performs more fine-grained segmentation of the scene and maintains clear segmentation boundaries. For example, in the segmentation completion result of cars, we predict clear separation of each car. In contrast, other methods show continuous semantic errors for occluded cars. In addition, our flow can effectively deal with the problem of mutual occlusion between different objects. Finally, we provide a video in the appendix to show the performance more intuitively.
@@ Appendix C More Visualizations Results
Appendix D Discussions
Flowscene shows strong performance on the benchmark with an improved number of parameters. This is beneficial for deploying real-world autonomous driving applications. But the inference time of the model needs to be improved. Semantic scene completion in multi-camera settings is also worth attention, which is our future work. Meanwhile, the legal challenges of autonomous driving as well as privacy and data security risks are still topics of debate. Finally, the robustness of semantic scene completion is also an issue worth exploring.

### v2

@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
This work was supported in part by the State Key Laboratory of Advanced Design and Manufacturing Technology for Vehicle under Grant 72365001, in part by the National Natural Science Foundation of China(Grant Nos. 62225205, 62472162, 62473137), in part by the National Key Researchand Development Program of China (No.2025YFB3003601), in part by the Major Science and Technology Research Projects of Hunan Province (Grant Nos. 2024QK2010, 2024QK2009).
Technical Appendices and Supplementary Material
This technical appendices consists of the following sections:
  • In Section A, we provide information of datasets, metrics and detailed implementation details.
  • In Section B, we provide more experiments results and analysis.
@@
  • In Section C, we analyze the limitations of our approach, directions for future work, and the broader impacts of this work.
We also include a video in the supplementary material.
Appendix A Experimental Setup
Datasets.
The SemanticKITTI[behley2019semantickitti, Geiger2012kitti] dataset includes dense semantic scene completion annotations and labels a voxelized scene with 20 semantic classes. It consists of 10 training sequences, 1 validation sequence, and 11 testing sequences. RGB images are resized to $1280\times 384$ for input processing. The SSCBench-KITTI-360[li2023sscbench, Liao2022kitti360] dataset contains 7 training sequences, 1 validation sequence, and 1 testing sequence, covering 19 semantic classes in total. The RGB images are resized to $1408\times 384$ for input processing.
Metrics.
Training Details.
We use RepVit [wang2023repvit] and FPN [lin2017feature] to extract features for all images. The number of historical temporal frames n is set to 2. We use and freeze the GMFlow [xu2022gmflow] optical flow estimation model to obtain optical flow information. We use the LSS paradigm for 2D-3D projection. The neighborhood cross-attention range is set to 7, and the number of attention heads is set to 8. Finally, the final outputs of SemantiKITTI is 20 classes, and SSCBench-KITTI-360 is 19 classes. All datasets have the scene size of $51.2m\times 51.2m\times 64m$ with the voxel grid size of $256\times 256\times 32$. By default, the model is trained for 25 epochs. We optimise the process, utilizing the AdamW optimizer with an initial learning rate of 1e-4 and a weight decay of 0.01. We also employ a multi-step scheduler to reduce the learning rate. All models are trained on two A100 Nvidia GPUs with 80G memory and batch size 4.
Implementation of Flow Consistency Check.
1. Variable Definition:
Forward Flow: ( $Flow^{t\rightarrow t-1}$ )
  • Maps pixels from the current frame ( $I_{t}$) to the previous frame ( $I_{t-1}$ ).
  • For a pixel (${x}_{t}\in I_{t}$), the corresponding location in ($I\_{t-1}$) is:
    ${x}_{t-1}={x}_{t}+Flow^{t\rightarrow t-1}({x}_{t})$
Backward Flow: ($Flow^{t-1\rightarrow t}$ )
  • Maps pixels from the previous frame ($I_{t-1}$) to the current frame ( $I_{t}$ ).
  • For a pixel ( ${x}_{t-1}\in I\_{t-1}$), the corresponding location in ( $I_{t}$ ) is:
    ${x}_{t}^{\prime}={x}\_{t-1}+Flow^{t-1\rightarrow t}({x}\_{t-1})$
2. Consistency Check
The forward-backward consistency check verifies whether a pixel mapping is valid by ensuring round-trip correspondence.
Round-trip Mapping
A pixel in ( $I_{t}$) is mapped to ( $I_{t-1}$) using forward flow, and then mapped back using backward flow:
${x}_{t}^{\prime\prime}={x}_{t}+Flow^{t\rightarrow t-1}({x}_{t})+Flow^{t-1\rightarrow t}({x}_{t}+F^{t\rightarrow t-1}({x}_{t}))$
Define the consistency residual as:
$\Delta({x}_{t})=Flow^{t\rightarrow t-1}({x}_{t})+Flow^{t-1\rightarrow t}({x}_{t}+F^{t\rightarrow t-1}({x}_{t}))$
If the norm ( $\|\Delta({x}_{t})\|$) is small (below a threshold), the mapping is considered consistent.
3. Occlusion Mask
Pixels with high inconsistency are typically considered occluded or unreliable.
Occlusion Mask (M(x)M({x})):
$M({x})=\begin{cases}1&\text{if }\|\Delta({x})\|>\tau\quad\text{(occluded)}\\ \\ 0&\text{otherwise}\quad\text{(non-occluded)}\end{cases}$
Where ($\tau$) is a predefined threshold.
Implementation of FGTA and OGVR module.
Algorithms 1 and 2 describe the implementation details of the two key components proposed in this work: Flow-Guided Temporal Aggregation (FGTA) and Occlusion-Guided Voxel Refinement (OGVR).
Algorithm 1 outlines the inference procedure for FGTA. Given the current and historical image frames $\{I_{t-i}\}_{i=0}^{N}$ and their corresponding features $\{F_{t-i}\}_{i=0}^{N}$, the algorithm first estimates forward and backward optical flows between the current frame $I_{t}$ and each historical frame $I_{t-i}$. Each historical feature map $F_{t-i}$ is warped to the reference frame using flow-guided warping. Cosine similarity between the warped features and the current frame feature $F_{t}$ is computed to assign adaptive weights, which emphasize temporally consistent and visually similar regions. Simultaneously, an occlusion mask is constructed through a bidirectional flow consistency check to identify unreliable regions. The weighted historical features are aggregated into $F_{agg}$, and a Neighborhood Cross-Attention (NCA) operation is applied to further refine $F_{t}$, focusing only on non-occluded regions. This process enhances temporal consistency and robustness to motion and occlusion artifacts.
Algorithm 2 describes the inference procedure for the OGVR module. The goal is to refine the voxel features in 3D space by integrating information from the aggregated feature volume $V_{agg}$, the current frame’s voxel features $V_{t}$, and the occlusion mask volume $V_{mask}$. Non-occluded regions are updated using features from $V_{agg}$, while occluded regions are filled in using the current frame’s voxel features $V_{t}$, which are more reliable in such areas. A per-voxel weight map is maintained to track the number of valid sources contributing to each voxel. The final voxel feature representation $V_{fine}$is obtained by normalizing the fused features with the accumulated weights, ensuring smooth transitions at the boundaries between occluded and non-occluded regions. Importantly, this refinement process is lightweight and introduces no additional parameter overhead.
Together, these modules enable FlowScene to effectively align temporal information and reason across occlusions in both 2D and 3D spaces, leading to improved semantic and geometric scene understanding in dynamic environments.
Algorithm 1 Inference algorithm of flow-guided temporal alignment
1:Inputs: Images $\{I_{t-i}\}_{i=0}^{N}$, Features $\{F_{t-i}\}_{i=0}^{N}$
2:$M=Zeros$ $\triangleright$ init occlusion mask
3:for $i=1$ to NN do
4:  $Flow^{t\rightarrow t-i},Flow^{t-i\rightarrow t}\leftarrow\mathcal{F}(I_{t},I_{t-i})$ $\triangleright$ compute dual optical flow
5:  $F_{warp}^{t-i\rightarrow t}\leftarrow\mathcal{W}arp(F_{t-i},Flow^{t\to t-i})$ $\triangleright$ flow-guided warp
6:  $w_{t-i\rightarrow t}\leftarrow\operatorname{similarity}(F_{warp}^{t-i\rightarrow t},F_{t})$ $\triangleright$ compute similarity weight
7:  ${F}_{agg}[i]\leftarrow w_{t-i\rightarrow t}\cdot F^{t-i\rightarrow t}_{warp}$ $\triangleright$ get the features corresponding to the weights
8:  $M\leftarrow\mathcal{CC}(Flow^{t\rightarrow t-i},Flow^{t-i\rightarrow t})\cup M$ $\triangleright$ compute occlusion mask
9:end for
[…]
12:Outputs: $F_{agg},F_{t},M$
Algorithm 2 Inference algorithm of occlusion-guided voxel refinement
1:Inputs: Aggregated voxel feature $V_{agg}$, current voxel feature $V_{t}$, occlusion mask $V_{mask}$
2:Initialize $\text{weight}\leftarrow\mathbf{0}$ with shape as $V_{t}$
3:Initialize $V_{fine}\leftarrow\mathbf{0}$ with shape as $V_{t}$
[…]
13:Outputs: Refined voxel feature $V_{fine}$
Table 11: Reproduce SOTA method using different image encoder.
Method Backbone mIoU(%) Params(M)
MonoScene [cao2022monoscene] EfficientNetB7 12.96 132.4
RepViT 12.59 91.0
VoxFormer [li2023voxformer] ResNet50 13.35 57.9
RepViT 13.62 54.7
BRGScene [li2023stereoscene] EfficientNetB7 15.43 161.4
RepViT 16.13 120.0
HTCL [li2024htcl] EfficientNetB7 17.13 181.4
RepViT 16.86 140.0
VLScene [wang2025vlscene] EfficientNetB7 17.44 88.8
RepViT 17.83 47.4
Ours EfficientNetB7 17.63 93.8
ResNet50 16.98 55.6
RepViT 18.13 52.4
Appendix B More Results
B.1 Reproduce SOTA Method using Different Image Encoder
To ensure a fair and consistent comparison across methods, we re-implemented several state-of-the-art Semantic Scene Completion models using a unified experimental setup. Specifically, we replaced the original image encoders used in existing methods with a lightweight and efficient backbone—RepViT [wang2023repvit]—while keeping all other architectural and training settings unchanged.
As shown in Table 11, the results demonstrate that RepViT significantly reduces the number of parameters across all models, often without degrading performance—and in some cases, even improving it. This confirms the generalizability and efficiency of RepViT as an image encoder for SSC tasks.
B.2 Standard Errors and Standard Deviations Results
We conducted a statistical analysis on the SemanticKITTI validation set and report the weighted mean IoU (W-mIoU), weighted standard deviation (W-SD), and weighted standard error (W-SE) across semantic categories. As shown, our method not only achieves the highest mIoU and W-mIoU, but also demonstrates W-SD and W-SE compared to other strong baselines. This indicates that our performance improvements are statistically meaningful and stable across classes.
Table 12: Standard Errors and Standard Deviations Results
Method mIoU($\uparrow)$ W-mIoU($\uparrow$) W-SD($\downarrow$) W-SE($\downarrow$)
Ours 17.70 33.11 13.80 6.50
HTCL (ECCV'2024) 17.08 32.64 14.17 6.67
CGFormer (NIPS'2024) 16.63 31.91 14.43 6.80
BRGScene (IJCAI'2024) 15.35 30.22 14.01 6.60
TPVFormer (ICCV'2023) 12.32 24.44 14.34 6.75
MonoScene (CVPR'2022) 11.08 22.58 14.38 6.77
B.3 More Quantitative Results
@@
To provide a more thorough comparison, we provide additional quantitative results of semantic scene completion on the SemanticKITTI validation set in Table 13. The results further demonstrate the effectiveness of our approach in enhancing 3D scene perception performance. Compared with the previous state-of-the-art methods, FlowScene is superior to other HTCL [li2024htcl] in semantic scene understanding, with a 1.00% increase in mIoU. In addition, compared with Symphonize [jiang2024symphonize], huge improvements are made in both occupancy and semantics. IoU and mIoU enhancement are of great significance for practical applications. It proves that we are not simply reducing a certain metric to achieve semantic scene completion.
Table 13: Quantitative results on the SemanticKITTI validation set. The best and the second best results are in bold and underlined, respectively.
@@
MonoScene [cao2022monoscene] CVPR’2022 S 36.86 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25 11.08
TPVFormer [huang2023tri] CVPR’2023 S 35.61 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52 11.36
OccFormer[zhang2023occformer] ICCV’2023 S 36.50 58.85 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.62 2.78 2.82 0.00 5.61 4.26 2.86 13.46
Symphonize [jiang2024symphonize] CVPR’2024 S 41.92 56.37 27.58 15.28 0.95 21.64 28.68 20.44 2.54 2.82 13.89 25.72 6.60 30.87 3.52 2.24 0.00 8.40 9.57 5.76 14.89
VoxFormer-T[li2023voxformer] CVPR’2023 T 44.15 53.57 26.52 19.69 0.42 19.54 26.54 7.26 1.28 0.56 7.81 26.10 6.10 33.06 1.93 1.97 0.00 7.31 9.15 4.94 13.35
H2GFormer [wang2024h2gformer] AAAI’2024 T 44.69 57.00 29.37 21.74 0.34 20.51 28.21 6.80 0.95 0.91 9.32 27.44 7.80 36.26 1.15 0.10 0.00 7.98 9.88 5.81 14.29
HASSC [wang2024HASSC] CVPR’2024 T 44.58 55.30 29.60 25.90 11.30 23.10 23.00 2.90 1.90 1.50 4.90 24.80 9.80 26.50 1.40 3.00 0.00 14.30 7.00 7.10 14.74
HTCL [li2024htcl] ECCV’2024 T 45.51 63.70 32.48 23.27 0.14 24.13 34.30 20.72 3.99 2.80 11.99 26.96 8.79 37.73 2.56 2.70 0.00 11.22 11.49 6.95 17.13
@@
B.4 Failure Case
We provide two failure cases in Figure 8.
Figure 8: Failure cases.
B.5 More Visualizations Results
We show visualization examples on the Semantickitti validation set, as shown in Figure 9. From left to right are the input image, the corresponding optical flow and occlusion mask, the front view SSC, and the top view SSC. Due to the motion information brought by the optical flow, the location information of the scene objects is more accurate and the layout is more reasonable. We report the performance of more visual comparison results on the SemanticKITTI validation set in Figure 10. We compare with VoxFormer [li2023voxformer] and BRGScene [li2023stereoscene]. In general, our method performs more fine-grained segmentation of the scene and maintains clear segmentation boundaries. For example, in the segmentation completion result of cars, we predict clear separation of each car. In contrast, other methods show continuous semantic errors for occluded cars. In addition, our flow can effectively deal with the problem of mutual occlusion between different objects. Finally, we provide a video in the appendix to show the performance more intuitively.